# Rethinking the Capacity of Graph Neural Networks for Branching Strategy

### v2

@@ 5. Numerical results
Experimental settings: We train an MP-GNN and a 2-FGNN with $L=2$, where we replace the functions $f^{l}(t_{j}^{l-1},A_{ij})$ and $g^{l}(s_{i}^{l-1},A_{ij})$ in the MP-GNN by $A_{ij}f^{l}(t_{j}^{l-1})$ and $A_{ij}g^{l}(s_{i}^{l-1})$ to guarantee that they take the value 00 whenever $A_{ij}=0$. For both GNNs, $p^{0},q^{0}$ are parameterized as linear transformations followed by a non-linear activation function; $\{p^{l},q^{l},f^{l},g^{l}\}_{l=1}^{L}$ are parameterized as 3-layer multi-layer perceptrons (MLPs) with respective learnable parameters; and the output mapping $r$ is parameterized as a 2-layer MLP. All layers map their input to a 1024-dimensional vector and use the ReLU activation function. With $\theta$ denoting the set of all learnable parameters of a network, we train both MP-GNN and 2-FGNN to fit the SB scores of the MILP dataset $\mathcal{G}$, by minimizing $\frac{1}{2}\sum_{G\in\mathcal{G}}\|F_{\theta}(G)-\textup{SB}(G)\|^{2}$ with respect to $\theta$, using Adam [kingma2014adam]. The networks and training scheme is implemented with Python and TensorFlow [abadi2016tensorflow]. The numerical experiments are conducted on a single NVIDIA Tesla V100 GPU for two datasets:
  • We randomly generate $100$ MILP instances, with 6 constraints and 20 variables, that are MP-tractable with probability $1$. SB scores are collected using SCIP [scip, scip2]. More details about instance generation are provided in Appendix D.
@@
Experimental results: The numerical results are displayed in Figure 3. One can see from Figure 3(a) that both MP-GNN and 2-FGNN can approximate the SB scores over the dataset of random MILP instances very well, which validates Theorem 4.4 and Theorem 4.7. As illustrated in Figure 3(b), 2-FGNN can perfectly fit the SB scores of (4.1) and (4.2) simultaneously while MP-GNN can not, which is consistent with Theorem 4.5 and Theorem 4.7 and serves as a numerical verification of the capacity differences between MP-GNN and 2-FGNN for SB prediction. The detailed exploration of training and performance evaluations of GNNs is deferred to future work to maintain a focused investigation on the theoretical capabilities of GNNs in this paper.
(a) Both MP-GNN and 2-FGNN can fit the SB scores of MP-tractable MILPs.
(b) 2-FGNN can fit SB scores of MP-intractable MILPs (4.1) and (4.2) while MP-GNN can not.
Figure 3. Numerical results of MP-GNN and 2-FGNN for SB score fitting
@@ 6. Conclusion
In this work, we study the expressive power of two types of GNNs for representing SB scores. We find that MP-GNNs can accurately predict SB results for MILPs within a specific class termed “message-passing-tractable” (MP-tractable). However, their performance is limited outside this class. In contrast, 2-FGNNs, which update node-pair features instead of node features as in MP-GNNs, can universally approximate the SB scores on every MILP dataset or for every MILP distribution. These findings offer insights into the suitability of different GNN architectures for varying MILP datasets, particularly considering the ease of assessing MP-tractability.
We also comment on limitations and future research topics. Although the universal approximation result is established for MP-GNNs and 2-FGNNs to represent SB scores, it is still unclear what is the required complexity/number of parameters to achieve a given precision. It would thus be interesting and more practically useful to derive some quantitative results. In addition, our current numerical experiments are somehow of small-scale. Implementing large-scale experiments and exploring efficient training strategies could lead to more practical impacts.
@@ Acknowledgements
Theorem A.2 ([chen2022representing-lp]*Theorem A.2).
@@
For any $G\sim\bar{G}\in\mathcal{G}_{m,n}^{\textup{MP}}$ with $\textup{SB}(G)\in\mathbb{R}^{n}$ and $\textup{SB}(\bar{G})\in\mathbb{R}^{n}$, the followings are true:
@@
Theorem A.4 (Lusin’s theorem [evans2018measure]*Theorem 1.14).
@@
Theorem A.5 ([chen2022representing-lp]*Theorem E.1).
@@
    For any $\sigma_{V}\in S_{m},\sigma_{W}\in S_{n}$, and $G\in X$, it holds that $\Phi((\sigma_{V},\sigma_{W})\ast G)=\sigma_{W}(\Phi(G))$.
@@
It is clear that $G\stackrel{{\scriptstyle W}}{{\sim}}_{2}\sigma_{W}\ast\bar{G}$ implies that $G\sim_{2}\bar{G}$. We then prove the reverse direction, i.e., $G\sim_{2}\bar{G}$ implies $G\stackrel{{\scriptstyle W}}{{\sim}}_{2}\sigma_{W}\ast\bar{G}$ for some $\sigma_{W}\in S_{n}$. It suffices to consider $L$ and hash functions such that there are no collisons in Algorithm 2 and no strict color refinement in the $L$-th iteration when $G$ and $\bar{G}$ are the input, which means that two edges are assigned with the same color in the $L$-th iteration if and only if their colors are the same in the $(L-1)$-th iteration. For any $j_{1},j_{2},j_{1}^{\prime},j_{2}^{\prime}\in W$, it holds that
@@ C.2. SB scores of MILPs distinguishable by 2-FWL test
Choose $L$ and hash functions such that there are no collisons in Algorithm 2 and no strict color refinement in the $L$-th iteration when $G$ and $\bar{G}$ are the input. Fix any $j\in W$ and construct the partions =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ as follows:
@@ C.3. Equivalence between the separation powers of the 2-FWL test and 2-FGNNs
Consider any $F\in\mathcal{F}_{\textup{2-FGNN}}$ with $L$ layers and let $s_{ij}^{l},t_{j_{1}j_{2}}^{l}$ and $\bar{s}_{ij}^{l},\bar{t}_{j_{1}j_{2}}^{l}$ be the features in the $l$-th layer of $F$. Choose $L$ and hash functions such that there are no collisons in Algorithm 2 when $G$ and $\bar{G}$ are the input. We will prove the followings by induction for $0\leq l\leq L$:
@@
Appendix D Details about numerical experiments
@@
We implement MP-GNN and 2-FGNN with Python 3.6 and TensorFlow 1.15.0 [abadi2016tensorflow]. Our implementation is built by extending the MP-GNN implementation in [gasse2019exact] [^5] See https://github.com/ds4dm/learn2branch.. The SB scores of randomly generated MILP instances are collected using SCIP [scip].

### v3

@@ 5. Numerical results
Experimental settings
We train an MP-GNN and a 2-FGNN with $L=2$, where we replace the functions $f^{l}(t_{j}^{l-1},A_{ij})$ and $g^{l}(s_{i}^{l-1},A_{ij})$ in the MP-GNN by $A_{ij}f^{l}(t_{j}^{l-1})$ and $A_{ij}g^{l}(s_{i}^{l-1})$ to guarantee that they take the value 00 whenever $A_{ij}=0$. For both GNNs, $p^{0},q^{0}$ are parameterized as linear transformations followed by a non-linear activation function; $\{p^{l},q^{l},f^{l},g^{l}\}_{l=1}^{L}$ are parameterized as 3-layer multi-layer perceptrons (MLPs) with respective learnable parameters; and the output mapping $r$ is parameterized as a 2-layer MLP. All layers map their input to a 1024-dimensional vector and use the ReLU activation function. With $\theta$ denoting the set of all learnable parameters of a network, we train both MP-GNN and 2-FGNN to fit the SB scores of the MILP dataset $\mathcal{G}$, by minimizing $\frac{1}{2}\sum_{G\in\mathcal{G}}\|F_{\theta}(G)-\textup{SB}(G)\|^{2}$ with respect to $\theta$, using Adam [kingma2014adam]. The networks and training scheme is implemented with Python and TensorFlow [abadi2016tensorflow]. The numerical experiments are conducted on a single NVIDIA Tesla V100 GPU for two datasets:
  • We randomly generate $100$ MILP instances, with 6 constraints and 20 variables, that are MP-tractable with probability $1$. SB scores are collected using SCIP [scip]. More details about instance generation are provided in Appendix E.
@@
Experimental results
The numerical results are displayed in Figure 3. One can see from Figure 3(a) that both MP-GNN and 2-FGNN can approximate the SB scores over the dataset of random MILP instances very well, which validates Theorem 4.4 and Theorem 4.7. As illustrated in Figure 3(b), 2-FGNN can perfectly fit the SB scores of (4.1) and (4.2) simultaneously while MP-GNN can not, which is consistent with Theorem 4.5 and Theorem 4.7 and serves as a numerical verification of the capacity differences between MP-GNN and 2-FGNN for SB prediction. The detailed exploration of training and performance evaluations of GNNs is deferred to future work to maintain a focused investigation on the theoretical capabilities of GNNs in this paper.
(a) MP-tractable MILPs: Both MP-GNN and 2-FGNN can fit the SB scores.
(b) MP-intractable MILPs (4.1) and (4.2): 2-FGNN can fit SB scores while MP-GNN can not.
Figure 3. Numerical results of MP-GNN and 2-FGNN for SB score fitting. In the right figure, the training error of MP-GNN on MP-intractable examples does not decrease after however many epochs.
Number of parameters
In Figure 3(b), the behavior of MP-GNN remains unchanged regardless of the number of parameters used, as guaranteed by Theorem 4.5. This error is intrinsically due to the structure of MP-intractable MILPs and cannot be reduced by adding parameters. Conversely, 2-FGNN can achieve near-zero loss with sufficient parameters, as guaranteed by Theorem 4.7 and confirmed by our numerical experiments. To further verify this, we tested 2-FGNN with embedding sizes from 64 to 2,048. All models reached near-zero errors, though epoch counts varied, as shown in Table 1. The results suggest that larger embeddings improve model capacity to fit counterexamples. The gains level off beyond an embedding size of 1,024 due to increased training complexity.
Table 1. Epochs required to reach specified errors with varying embedding sizes for 2-FGNN.
Embedding size 64 128 256 512 1,024 2,048
Epochs to reach $10^{-6}$ error 16,570 5,414 2,736 1,442 980 1,126
Epochs to reach $10^{-12}$ error 18,762 7,474 4,412 2,484 1,128 1,174
Larger instances
While our study primarily focuses on theory and numerous empirical studies have shown the effectiveness of GNNs in branching strategies (as noted in Section 1), we conducted experiments on larger instances to further assess the scalability of this approach. We trained an MP-GNN on 100 large-scale set covering problems, each with 1,000 variables and 2,000 constraints, generated following the methodology in [gasse2019exact]. The MP-GNN achieved a training loss of $1.94\times 10^{-4}$, calculated as the average $\ell_{2}$ norm of errors across all training instances.
@@ 6. Conclusion
In this work, we study the expressive power of two types of GNNs for representing SB scores. We find that MP-GNNs can accurately predict SB results for MILPs within a specific class termed “message-passing-tractable” (MP-tractable). However, their performance is limited outside this class. In contrast, 2-FGNNs, which update node-pair features instead of node features as in MP-GNNs, can universally approximate the SB scores on every MILP dataset or for every MILP distribution. These findings offer insights into the suitability of different GNN architectures for varying MILP datasets, particularly considering the ease of assessing MP-tractability. We also comment on limitations and future research topics. Although the universal approximation result is established for MP-GNNs and 2-FGNNs to represent SB scores, it is still unclear what is the required complexity/number of parameters to achieve a given precision. It would thus be interesting and more practically useful to derive some quantitative results. In addition, exploring efficient training strategies or alternatives of higher order GNNs for MILP tasks is an interesting and significant future direction.
@@ Acknowledgements
Theorem A.2 ([chen2022representing-lp, Theorem A.2]).
@@
For any $G,\bar{G}\in\mathcal{G}_{m,n}^{\textup{MP}}$ with $\textup{SB}(G)\in\mathbb{R}^{n}$ and $\textup{SB}(\bar{G})\in\mathbb{R}^{n}$, the followings are true:
@@
Theorem A.4 (Lusin’s theorem [evans2018measure, Theorem 1.14]).
@@
Theorem A.5 ([chen2022representing-lp, Theorem E.1]).
@@
    For any $\sigma_{V}\in S_{m},\sigma_{W}\in S_{n}$, and $G\in X$, it holds that $\Phi((\sigma_{V},\sigma_{W})\ast G)=\sigma_{W}(\Phi(G))$, where $(\sigma_{V},\sigma_{W})\ast G$ represents the MILP-graph obtained from $G$ by reordering vertices with permutations $\sigma_{V}$ and $\sigma_{W}$.
@@
It is clear that $G\stackrel{{\scriptstyle W}}{{\sim}}_{2}\sigma_{W}\ast\bar{G}$ implies that $G\sim_{2}\bar{G}$. We then prove the reverse direction, i.e., $G\sim_{2}\bar{G}$ implies $G\stackrel{{\scriptstyle W}}{{\sim}}_{2}\sigma_{W}\ast\bar{G}$ for some $\sigma_{W}\in S_{n}$. It suffices to consider $L$ and hash functions such that there are no collisions in Algorithm 2 and no strict color refinement in the $L$-th iteration when $G$ and $\bar{G}$ are the input, which means that two edges are assigned with the same color in the $L$-th iteration if and only if their colors are the same in the $(L-1)$-th iteration. For any $j_{1},j_{2},j_{1}^{\prime},j_{2}^{\prime}\in W$, it holds that
@@ C.2. SB scores of MILPs distinguishable by 2-FWL test
Choose $L$ and hash functions such that there are no collisions in Algorithm 2 and no strict color refinement in the $L$-th iteration when $G$ and $\bar{G}$ are the input. Fix any $j\in W$ and construct the partions =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ as follows:
@@ C.3. Equivalence between the separation powers of the 2-FWL test and 2-FGNNs
Consider any $F\in\mathcal{F}_{\textup{2-FGNN}}$ with $L$ layers and let $s_{ij}^{l},t_{j_{1}j_{2}}^{l}$ and $\bar{s}_{ij}^{l},\bar{t}_{j_{1}j_{2}}^{l}$ be the features in the $l$-th layer of $F$. Choose $L$ and hash functions such that there are no collisions in Algorithm 2 when $G$ and $\bar{G}$ are the input. We will prove the followings by induction for $0\leq l\leq L$:
@@
Appendix D Extensions of the theoretical results
This section will explore some extensions of our theoretical results.
D.1. Extension to other types of SB scores
The same analysis for Theorem 4.4 and Theorem 4.7 still works as long as the SB score is a function of $f_{\textup{LP}}^{*}(G,j,l_{j},\hat{u}_{j})$, $f_{\textup{LP}}^{*}(G,j,\hat{l}_{j},u_{j})$, and $f_{\textup{LP}}^{*}(G)$:
  • We prove in Theorem A.3 that if two MILP-graphs are indistinguishable by the WL test, then they must be isomorphic and hence have identical SB scores (no matter how we define the SB scores). So Theorem 4.4 is still true.
  • We prove in Theorem C.4 that if two MILP-graphs are indistinguishable by 2-FWL test, then they have the same value of $f_{\textup{LP}}^{*}(G,j,l_{j},\hat{u}_{j})$ (and $f_{\textup{LP}}^{*}(G,j,\hat{l}_{j},u_{j})$). Therefore, Theorem C.3 still holds if the SB score is a function of $f_{\textup{LP}}^{*}(G,j,l_{j},\hat{u}_{j})$, $f_{\textup{LP}}^{*}(G,j,\hat{l}_{j},u_{j})$, and $f_{\textup{LP}}^{*}(G)$, which implies that Theorem 4.7 is still true.
Therefore, Theorems 4.4 and 4.7 work for both linear product score functions in [dey2024theoretical].
D.2. Extension to varying MILP sizes
While Theorems 4.4 and 4.7 assume MILP sizes $m$ and $n$ are fixed, we now discuss extending these results to data distributions with variable $m$ and $n$.
First, our theoretical results can be directly extended to MILP datasets or distributions where $m$ and $n$ vary but remain bounded. Following Lemma 36 in [azizian2020expressive], if a universal-approximation theorem applies to $\mathcal{G}_{m,n}$ for any fixed $m$ and $n$ (as shown in our work) and at least one GNN can distinguish graphs of different sizes, then the result holds across a disjoint union of finitely many $\mathcal{G}_{m,n}$.
If the distribution has unbounded $m$ or $n$, for any $\epsilon>0$, one can always remove a portion of the tail to ensure boundedness in $m$ and $n$. In particular, there always exist large enough $m_{0}$ and $n_{0}$ such that $\mathbb{P}[m(G)\leq m_{0}]\geq 1-\epsilon$ and $\mathbb{P}[n(G)\leq n_{0}]\geq 1-\epsilon$. The key point is that for any $\epsilon>0$, such $m_{0}$ and $n_{0}$ can always be found. Although these values may be large and dependent on $\epsilon$, they are still finite. This allows us to apply the results for the bounded-support case.
Note that the “tail removal” technique mentioned above comes from the fact that a probability distribution has a total mass of 1:
$1=\sum_{n=0}^{\infty}\mathbb{P}[n(G)=n]=\lim_{n_{0}\to\infty}\sum_{n=0}^{n_{0}}\mathbb{P}[n(G)=n]=\lim_{n_{0}\to\infty}\mathbb{P}[n(G)\leq n_{0}].$
By the definition of a limit, this clearly implies that for any $\epsilon>0$, there exists a sufficiently large $n_{0}$ such that $\mathbb{P}[n(G)\leq n_{0}]\geq 1-\epsilon$. A similar argument applies to $m$.
Appendix E Details about numerical experiments
@@
We implement MP-GNN and 2-FGNN with Python 3.6 and TensorFlow 1.15.0 [abadi2016tensorflow]. Our implementation is built by extending the MP-GNN implementation of [gasse2019exact] in https://github.com/ds4dm/learn2branch. The SB scores of randomly generated MILP instances are collected using SCIP [scip].